Find two numbers such that four less than twice the first number is two more than the second number, and the sum of the two numbers is nine. What are the two numbers. Solve and show all work.
Answer by checkley71(8403)

X+Y=9 & 2X-4=Y+2 OR X=9-Y THUS 2(9-Y)-4=Y+2 OR 18-2Y-4=Y+2 OR 14-2Y=Y+2 OR
14-2=Y+2Y OR 12=3Y OR Y=12/3 OR Y=4 THEN X+4=9OR X=9-4 OR X=5
PROOF 2*5-4=4+2 OR 10-4=6 OR 6=6